Cathy, the manager of Cathys Catering, Inc., uses activity-based costing to compute the costs of her catered parties. Each pRequired A Required Required C Compute the cost of a 20-guest evening formal dinner. Evening formal dinnerRequired A Required B Required How much should Cathy charge for each guest for each type of party if she wants to cover her c
Answer #1
a. Cost of afternoon picnic:
$
Advertising 100
Planning 75
Renting equipment [50+(10*20)] 250
Obtaining insurance 200
Serving (50*4) 200
Preparing food (20*20) 400
Total 1225
b. Cost of formal dinner:
$
Advertising 100
Planning 125
Renting equipment [75+(20*20)] 475
Obtaining insurance 400
Serving (75*4) 300
Preparing food (30*20) 600
Total 2000
c. Charging amount=Total cost/Number of guests
Charging amount:
Afternoon picnic=1225/20=$ 61.25 per guest
Formal dinner=2000/20=$ 100 per guest